OpenStudy (anonymous):

3-2i/-i please solve this is urgent.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

\[\frac{ 3-2i }{-i }*\frac{ i }{ i }\] =\[\frac{ 3i-2i ^{2} }{ -i ^{2} }\]

OpenStudy (anonymous):

It has to be In standard complex form

OpenStudy (anonymous):

5i

OpenStudy (anonymous):

\[i ^{2}=-1\] so \[3i+2\] is the answer

OpenStudy (anonymous):

so it is in form of ai+b

OpenStudy (anonymous):

Okay thanks.